OpenStudy (anonymous):

Given that f(x) = 4x − 3 and g(x) = quantity of 2x minus 1, divided by 3, solve for g(f(2)).

OpenStudy (anonymous):

1 3 9 15

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Sorry I forgot this stuff.

OpenStudy (anonymous):

what about The function p(x) = 2x +1 determines how many pizzas need to be purchased for an after school meeting, where x is the number of students at the meeting. The club leader uses r(p(x)) to find the amount of money to bring for the pizza purchase. The function r(x) = 4x − 3. Solve for how much money to bring when there are 9 students in the meeting. @iGreen

OpenStudy (igreen):

p(x) = 2(9) + 1 p(x) = 18 + 1 p(x) = 19

OpenStudy (igreen):